The production possibilities frontier of an economy that devotes a certain amount of resources to produce wheat and textiles is shown below.

PPF

If the economy moves along the PPF from point K to point G to point H, the opportunity cost of _____, and if it moves from H to point G to point K, the opportunity cost of _____.

Question 24 options:

1 sq. yard of textiles decreases; 1 bushel of wheat decreases

1 bushel of wheat increases; 1 sq. yard of textiles increases

1 sq. yard of textiles increases; 1 bushel of wheat decreases

1 bushel of wheat decreases; 1 sq. yard of textiles increases

1 sq. yard of textiles increases; 1 bushel of wheat increases
Answer #1

Opportunity cost of moving from K to G or of producing 1 sq yard of textile = (230-200)/(180-140)= 0.75

Opportunity cost of moving from G to H= (200-100)/(250-180)

= 1.43

1 sq yard of textile increases; 1 bushel of wheat decreases.